Cavendish Convenience Stores purchases 1000 U.S. Road Maps a year for its inventory from its supplier, who offers pricing at quantity discounts. The cost for Cavendish to place an order is $42, and the annual carrying cost is 15%/year. What quantity should Cavendish order? The quantities and pricing from this supplier are shown in the following table: ORDER QUANTITY UNIT PRICE 0-149 $12.00 150-349 $11.70 350-999 $11.50 1000 or more $11.20 a) What is the EOQ when the unit price is $12? b) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where the unit price is $11.70? c) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where the unit price is $11.50? d) What is the total cost (holding + purchase+ ordering) when purchasing from the supplier where the unit price is $11.20? e) Based on the total costs, what quantity should be ordered from the supplier?
Answer #1

Please find below the response to part a -d in the table below.

Order quantity of 350 yields lowest total cost (holding + purchase+ ordering) and must be ordered.

Annual demand, D 1000
Order cost, Co $42
Quantity 0-149 150-349 350-999 1000+
Cost, C $12.00 $11.70 $11.50 $11.20
Holding cost , Cc $1.80 $1.76 $1.73 $1.68
(15%*12) (15%*11.70) (15%*11.50) (15%*11.20)
EOQ = sqrt(2DCo/Cc) 216.02 218.78 220.67 223.61
Considered Order quantity - Q for calculation 149.00 218.78 350 1000
Considered 149, max order range as EOQ is above the range Q= 218.78 , EOQ Consider min 350 to get a price discount as EOQ is below the range Consider min 1000 to get a price discount
Order size 0-149 150-349 350-999 1000+
Road maps cost (Purchase)= D*Cost, Different Costs $12,000.0 $11,700.0 $11,500.0 $11,200.0
Holding cost = (Q/2) * Cc, Differenct Q as mentioned above $134.10 $191.98 $301.88 $840.00
Ordering cost = (D/Q) * Co $281.88 $191.97 $120.00 $42.00
Total cost = Sum of above 3 $12,415.98 $12,083.95 $11,921.88 $12,082.00